how do i find the height of a cylinder with a surface area of 164cm squared, with a radius of 4cm.
Answer by venugopalramana(3286)

how do i find the height of a cylinder with a surface area of 164cm squared, with a radius of 4cm.
it is assumed that the surface area mentioned is lateral surface area and not total suraface area including top ,bottom and lateral surface area.
lateral surface area of a cylinder =2(pie)rh=164..taking pie as approximately equal to 22/7
2*(22/7)*4h=164
h=164*7/(2*22*4)=287/44...or =41/2pie